Please help me solve this system and graph 6u+v=32, 6u=v+40
Answer by Alan3354(69443)

Please help me solve this system and graph 6u+v=32, 6u=v+40
--------------
6u+v=32
6u=v+40
--------------
6u+v=32
6u-v=40
------- Add
===================
Both eqns are linear, straight lines.
You need 2 points to plot a straight line.
Pick 2 values for 1 variable and find the other.
Plot the 2 points.
Draw a line thru them.
========================
Do the same for the other eqn.
Notice that the intersection is the solution to the system.